4. Prepare an unadjusted trial balance.

Required information

[The following information applies to the questions displayed below.]
    

The general ledger of Jackrabbit Rentals at January 1, 2021, includes the following account balances:

Accounts Debits Credits
Cash $ 48,500
Accounts Receivable 32,700
Land 117,800
Accounts Payable 16,000
Notes Payable (due in 2 years) 37,000
Common Stock 107,000
Retained Earnings 39,000
Totals $ 199,000 $ 199,000

The following is a summary of the transactions for the year:

1. January 12 Provide services to customers on account, $69,400.
2. February 25 Provide services to customers for cash, $78,800.
3. March 19 Collect on accounts receivable, $46,400.
4. April 30 Issue shares of common stock in exchange for $37,000 cash.
5. June 16 Purchase supplies on account, $13,500.
6. July 7 Pay on accounts payable, $12,000.
7. September 30 Pay salaries for employee work in the current year, $71,200.
8. November 22 Pay advertising for the current year, $23,200.
9. December 30 Pay $3,600 cash dividends to stockholders.

The following information is available for the adjusting entries.

Accrued interest on the notes payable at year-end amounted to $3,200 and will be paid January 1, 2022. Accrued salaries at year-end amounted to $2,200 and will be paid on January 5, 2022. Supplies remaining on hand at the end of the year equal $3,000.

4. Prepare an unadjusted trial balance.
Answer #1

we will journalize the entries for easy understanding

January 12 account receivable 69400
service revenue 69400
Feb 25 cash 78800
service revenue 78800
MAR 19 cash 46400
account receivable 46400
Apr 30 cash 37000
common stock 37000
June 16 supplies 13500
account payable 13500
July 7 account payable 12000
cash 12000
Sept 30 salary 71200
cash 71200
Nov 22 advertising expense 23200
cash 23200
December 30 dividend 3600
cash 3600

find closing balance before adjustment

cash = 48500+78800+46400+37000-12000-71200-23200-3600

=100700

account receivable = 32700+69400-46400 = 55700

account payable =16000+13500-12000=17500

common stock = 37000+107000=144000

service revenue=69400+78800=148200

unadjusted trial balance

debit credit
cash 100700
Account receivable 55700
supplies 13500
land 117800
account payable 17500
Notes payable 37000
common stock 144000
dividend 3600
retained earnings 39000
service revenue 148200
advertising expense 23200
salary expense 71200